# Autocovariance and Optimal Design for Random Walk Metropolis-Hastings Algorithm

### v1

@@ IV Optimal Proposal Design
Since $h(\ell)$ coincides with the efficiency criterion in the diffusion limit analysis of Gelman and Roberts, the optimizer $\ell^{*}\mathrel{\mathop{\ordinarycolon}}=\arg\max_{\ell}h(\ell)=2.38/\sqrt{\bar{m}}$ yields the well-known optimal acceptance rate $\mathbb{E}[\alpha(x,y)]\approx 0.23$, see [RobertsandGelman], while simultaneously minimizing the diagonal entries of the unit-lag covariance matrix by Theorem 4.3.
@@ V Numerical Study
For the proposal distribution, we consider two cases: a Gaussian–component symmetric bimodal density $\phi(x)=\tfrac{1}{2}\mathcal{N}(x^{*},\sigma_{q}^{2})+\tfrac{1}{2}\mathcal{N}(-x^{*},\sigma_{q}^{2})$ with $\sigma_{q}>0$, and a standard Gaussian symmetric unimodal density. To ensure a fair comparison, we evaluate each MH chain in terms of effective sample size (ESS) and unit-lag autocorrelation, matched at the same (or nearly the same) average acceptance rate and running time (In practice, we fixed the wall-clock budge). The ESS is defined as
@@
where $N$ is the total number of samples and $\rho_{k}$ is the lag-$k$ autocorrelation [DefofESS]. ESS measures the number of independent samples that the correlated MCMC output is equivalent to, thereby quantifying the impact of autocorrelation on estimation efficiency. Across target densities, the bimodal design typically yields higher ESS and lower unit-lag autocorrelation than the standard Gaussian proposal.
@@ VI Concluding Remarks and Future Work
Understanding the autocovariance structure of the Metropolis–Hastings algorithm is essential for assessing its efficiency. In this paper, we presented new theoretical results for random-walk MH chains targeting symmetric unimodal densities and introduced an optimal proposal design based on the unit-lag covariance, ultimately relating the analysis to high-dimensional settings in an asymmetric sense. Future work will focus on relaxing the restrictive assumptions on target distributions and extending the framework to multi-lag covariance structures.
@@ Acknowledgment
The author thanks Dr. James A. Fill for an earlier proof of Lemma 3.3 and for proving the equation (9), which plays a crucial role in establishing Theorem 3.5.

### v2

@@ IV Optimal Proposal Design
Since $h(\ell)$ coincides with the efficiency criterion in the diffusion limit analysis of Roberts et al., the optimizer $\ell^{*}\mathrel{\mathop{\ordinarycolon}}=\arg\max_{\ell}h(\ell)=2.38/\sqrt{\bar{m}}$ yields the well-known optimal acceptance rate $\mathbb{E}[\alpha(\boldsymbol{X},\boldsymbol{Y})]\approx 0.23$, see [13], while simultaneously minimizing the diagonal entries of the unit-lag covariance matrix by Theorem 4.3.
@@ V Numerical Study
For the proposal distribution, we consider two cases: a Gaussian–component symmetric bimodal density $\phi(x)=\tfrac{1}{2}\mathcal{N}(x^{*},\sigma_{q}^{2})+\tfrac{1}{2}\mathcal{N}(-x^{*},\sigma_{q}^{2})$ with $\sigma_{q}>0$, and a standard Gaussian symmetric unimodal density. To ensure a fair comparison, we evaluate each MH chain in terms of effective sample size (ESS) and unit-lag autocorrelation, matched at the same (or nearly the same) average acceptance rate and running time (In practice, we fixed the wall-clock budget). The ESS is defined as
@@
where $N$ is the total number of samples and $\rho_{k}$ is the lag-$k$ autocorrelation [8]. ESS measures the number of independent samples that the correlated MCMC output is equivalent to, thereby quantifying the impact of autocorrelation on estimation efficiency. Across target densities, the bimodal design typically yields higher ESS and lower unit-lag autocorrelation than the standard Gaussian proposal.
To ensure a fair comparison across different average acceptance rates, we applied a global scaling factor to the proposal distributions. Specifically, for the bimodal proposal, the ratio $\sigma_{q}/x^{*}$ was held fixed to preserve the bimodal geometry, while a scalar tuning parameter was adjusted numerically to achieve the target average acceptance rate. This ensures that the performance differences observed in Figure 1 are attributable to the distributional shape rather than suboptimal variance scaling.
@@ VI Concluding Remarks and Future Work
Understanding the autocovariance structure of the Metropolis–-Hastings algorithm is essential for assessing its efficiency. In this paper, we presented new theoretical results for random-walk MH chains targeting symmetric unimodal densities and introduced an optimal proposal design based on the unit-lag covariance, ultimately relating the analysis to high-dimensional settings in an asymmetric sense. Future work will focus on relaxing the restrictive assumptions on target distributions and extending the framework to multi-lag covariance structures.
@@ Acknowledgment
The author thanks Dr. James A. Fill for an earlier proof of Lemma 3.3 and for helping us to formulate the statements and proofs of the results in Section 3 (especially Theorems 3.4–3.5).